OpenStudy (anonymous):

4 + x/2 > 2x How do I solve it? Please and Thank you!

OpenStudy (anonymous):

4 + x/2 > 2x multiply both sides by 2 to get 4 + x > 4x bring x over 4> 4x - x 4 >3x 4/3 > x

hero (hero):

You multiplied both sides by 2 but still have 4 on the left side?

OpenStudy (anonymous):

i saw it as (4 + x)/2

hero (hero):

4 + x/2 > 2x 4 + x/2 >4x/2 4 > 4x/2 - x/2 4 >(4x - x)/2 4 > 3x/2 4*2 > 3x 8 > 3x 8/3 > x

hero (hero):

I don't know which one it is either

OpenStudy (anonymous):

well i guess the writer will have to verify

hero (hero):

@akb48, is it \[\frac{4 +x}{2} > 2x\] or \[4 + \frac{x}{2} > 2x\]

OpenStudy (anonymous):

The second one.

OpenStudy (anonymous):

multiply by 2 you get 8 + x > 4x subtracting x 8 > 3x dividing by 3 8/3 > x x < 8/3 \[x \in (-\infty, \frac{ 8 }{ 3 } )\]